Bellingham Company produces a product that requires 14 standard pounds per unit. The standard price is $11.5 per pound. If 3,100 units used 44,300 pounds, which were purchased at $11.15 per pound, what is the direct materials (a) price variance, (b) quantity variance, and (c) cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

a. Direct materials price variance $ Favorable
b. Direct materials quantity variance $ Unfavorable
c. Direct materials cost variance $    Favorable

2.

Bellingham Company produces a product that requires 8 standard direct labor hours per unit at a standard hourly rate of $17.00 per hour. If 3,400 units used 26,100 hours at an hourly rate of $17.85 per hour, what is the direct labor (a) rate variance, (b) time variance, and (c) cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

a. Direct labor rate variance $ Unfavorable
b. Direct labor time variance $ Favorable
c. Direct labor cost variance $ Unfavorable
Answer #1
1
a. Direct materials price variance -15505 Favorable =44300*(11.15-11.5)
b. Direct materials quantity variance 10350 Unfavorable =11.5*(44300-3100*14)
c. Direct materials cost variance -5155 Favorable =(44300*11.15)-(3100*14*11.5)
2
a. Direct labor rate variance 22185 Unfavorable =26100*(17.85-17)
b. Direct labor time variance -18700 Favorable =17*(26100-3400*8)
c. Direct labor cost variance 3485 Unfavorable =(26100*17.85)-(3400*8*17)